OpenStudy (anonymous):

Find an equation of the line that satisfies the given conditions: Through (-2,11); perpendicular to the line 4x-8y=1 Detailed steps if necessary, please and many thanks.

OpenStudy (anonymous):

solve for y for the equation given.. \[y = \frac{1}{2}x - \frac{1}{8}\] in order for a function to be perpendicular to another function the slopes must be negative reciprocals. So the negative reciprocal of 1/2 is -2. That is the slope for the new line. Now by using the point-slope formula you get the equation for the new line\[y - 11 = -2(x-(-2))\] Simplify and solve for y and you get \[y = -2x + 7\]